OpenStudy (loser66):

A-130kg tackler moving at 2.5m/s meets head on and tackles an 82kg halfback moving at 5.0m/s. What will be their mutual speed immediately after the collision? Please explain me.

OpenStudy (anonymous):

This sounds like a conservation of energy equation. You got the mass and initial velocity of both bodies. The question then asks you what's their mutual speed which in states that are are connected together as one body, both of them are experiencing the same velocity after collision. Your set up equation would look as so: \[\frac{ 1 }{ 2 }m _{1}v _{1}^{2}+\frac{ 1 }{ 2}m _{2}v _{2}^{2} = \frac{ 1 }{ 2 }(m _{1}+m _{2})v _{f}^{2}\]It's unknown which way the half back is running, so I'm just going to assume it has a positive velocity. Or if they are coming at each other, one of their velocities will have to be negative, since they are traveling in opposite directions. plug in your values, where m1 and m2 represents the mass values of the people respectfully and v1 and v2 is their respective initial velocities. You now solve for v_f, their final velocity after collision. Where you see (m1+m2) is their combined mass as they are acting as one object now.

OpenStudy (loser66):

what does 82kg halfback mean? I am sorry for dumb question. My English is not good to completely understand it

OpenStudy (loser66):